OpenStudy (anonymous):

If we observe a star over 6 months and obtain a parallax angle (p) for the star, how can we calculate the distance to the star? A. distance = p2 B. distance = 1/p C. distance = p-1

OpenStudy (amistre64):

catch a ride on a moon beam :)

OpenStudy (amistre64):

what defines a parallax angle?

OpenStudy (saifoo.khan):

http://www.astronomynotes.com/starprop/trig-anim.gif

OpenStudy (anonymous):

hmmmm

OpenStudy (anonymous):

moar pwease..kinna confused >:P

OpenStudy (phi):

let x = sun to earth distance = 1 au let d = distance to the star tan(p)= x/d or d= x/tan(p) for p<<1 (in radians) tan(p) ~ p so d = x/p in AU

OpenStudy (anonymous):

so i/p if you measure in parsecs